islayzombies115:

a hiker in Africa discovers a skull that contains 63 of its original amount of c-14. find the age of the skull to the nearest year

boyyy:

According to a formula that I found online, we have t = [ ln (N / N0) / (-0.693) ] * 5730 where t is the age of the fossil, N / N0 is the percent of C-14 that remains and 5730 is the half-life of C-14....so we have t = ( ln (.63) / (-0.693)) * 5730 = about 3820 yrs old

islayzombies115:

it says that is not the right answer

boyyy:

N / N0 = e^(-kt) 54% = e^(-kt) ln(.54) =-k t [ln(.54)] / (-0.0001) = t the half-life of ¹⁴C is 5730 yrs ___ since the skull contains more than half of the original amount, the age should be less than 5730 yr since the calculated age does not bear this out, it appears that the given value of k is incorrect using the half life ___ .54 = (1/2)^(t / 5730) [log(.54)] / [log(1/2)] = t / 5730 {[log(.54)] / [log(1/2)]} * 5730 = t

boyyy:

what about thiss

islayzombies115:

nope ;-;

boyyy:

whats the answer choices

islayzombies115:

its a type in answer box

boyyy:

try looking it up

islayzombies115:

a hiker in africa discovers a skull that contains 63 of its original amount of c-14. n = n0e-kt n0= intentional amount of c-14( at time t=0) n-= amount of c-14 at time t k=0.0001 t = time, in years find the age of the skull in the nearest year